OpenStudy (anonymous):

How do you use prime factorization to find the least common multiple of 63 and 49?

OpenStudy (anonymous):

just multiply 63*49

OpenStudy (anonymous):

but thats not using prime factorization. I have to use prime factorization

OpenStudy (anonymous):

3x3x7 and 7x7 Now what do I do?

OpenStudy (anonymous):

When you have that you just compare the two and check for common numbers. In this case 7 appears once in both. So 7 is the gcd

OpenStudy (anonymous):

I need the LCM though

OpenStudy (anonymous):

that is the same

OpenStudy (anonymous):

7 is not the least common multiple of 63 and 49, it's smaller than both of them

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

Oh I know, sorry. You are right That will be 3^2*7^2

OpenStudy (anonymous):

441

OpenStudy (anonymous):

You should count the distinct prime factors

OpenStudy (anonymous):

there is 7^2 and 7 so you just need the bigger

OpenStudy (anonymous):

where'd the 2 come from?

OpenStudy (anonymous):

49=7^2 63=3^2*7 LCM=49*63/7 because 7 is in both So LCM=3^2*7^2=441

OpenStudy (anonymous):

Ok thanks!